Hi I need the steps and answer confused a bit
Solve for x.
x - (2x - 6) = 3x - 2(x + 5)
: Hi I need the steps and answer confused a bit
Solve for x.
x - (2x - 6) = 3x - 2(x + 5)

Answer by stanbon(19008)

x - (2x - 6) = 3x - 2(x + 5)
x -2x + 6 = 3x -2x -10
-x + 6 = x -10
2x = 16
x = 8
=========
Cheers,
Stan H.